OpenStudy (babyslapmafro):

Find a horizontal line that divides the area between y=x^2 and y=9 into two equal parts.

OpenStudy (babyslapmafro):

I found the total area enclosed by the two curves (A=36) but how to a find a horizontal line that divides this area in half?

OpenStudy (anonymous):

you'd want to set up an integral such that its result would yield half of this area

OpenStudy (anonymous):

or in this case use symmetry. a parabola has equal area on either side of its line of symmetry. in this case the line of symmetry is x = 0

OpenStudy (anonymous):

oh i think i misread your question. you need a horizontal line. sorry about that! in this case you'd want to integrate with respect to y

OpenStudy (anonymous):

alternatively we could keep this with respect to x, and solve for the limits of integration: \[2\int\limits_{0}^{a} (9 - x^{2}) dx = 18\]

OpenStudy (anonymous):

\[\int\limits_{-a}^{a} (9 - x^{2}) dx = 18\] This is equivalent (not sure why the "-a" doesn't show up as the lower limit, but that is what was typed.)

OpenStudy (babyslapmafro):

Ok I tried to solve but I could not get the right answer. The answer in the book is \[\frac{ 9 }{ \sqrt[3]{4} }\]

OpenStudy (babyslapmafro):

I got two separate values, something like ~4.69 and ~1.08

OpenStudy (babyslapmafro):

@binarymimic

OpenStudy (anonymous):

Ah, only one of those answers is in the domain of the problem. sqrt(9 - x^2)

OpenStudy (babyslapmafro):

ok but both of my answer were off

OpenStudy (anonymous):

Sorry, 9 - x^2, not sqrt(9 - x^2)

OpenStudy (babyslapmafro):

\[18x-\frac{ 2x^3 }{ 3 }]^{a}_{0}=18\]

OpenStudy (babyslapmafro):

After that integrationi plugged into my calc and tried to find the intersection point to find a

OpenStudy (anonymous):

\[\int\limits_{-1.04189}^{1.04189} (9 - x^{2}) dx \approx 18\]

OpenStudy (anonymous):

should be a negative sign on the lower limit, but it doesn't show up anymore for some reason

OpenStudy (babyslapmafro):

answer in the book is \[\frac{ 9 }{ \sqrt[3]{4} }\]

OpenStudy (anonymous):

The book says the horizontal line that divides the area into two equal parts is: \[y = \frac{ 9 }{ \sqrt[3]{4} }\] ?

OpenStudy (babyslapmafro):

yes

OpenStudy (anonymous):

sorry openstudy is glitchy again